# Positivstellensätze for polynomial matrices with universal quantifiers

### v2

@@ 4. Positivstellensätze for polynomial matrices with UQs and without the Archimedean condition
The results in Section 3 are derived under the Archimedean condition on the quadratic module generated by $(G,\nu)$, which requires $\mathcal{X}$ to be bounded. Lots of efforts have been made to provide SOS-structured representations for polynomials that are PD on a unbounded semialgebraic set. The goal of this section is to extend some of these well-known results to the setting of polynomial matrices with UQs.
4.1. Putinar–Vasilescu-type Positivstellensatz
Let $\theta\coloneqq 1+\|{\boldsymbol{x}}\|^{2}$. For every polynomial $f\in{\mathbb{R}}[{\boldsymbol{x}}]$ nonnegative on a general basic semialgebraic set, Putinar and Vasilescu [42] proved that for a given $\varepsilon>0$ and $2d\geq\deg(f)$, there exists a nonnegative integer $k$ such that $\theta^{k}(f+\varepsilon\theta^{d})$ belongs to the quadratic module associated with the basic semialgebraic set. Using Jacobi’s technique [20], Mai et al. [34] provided an alternative proof of Putinar and Vasilescu’s result with an effective degree bound on polynomials involved in such certificates.
Building on Theorem 3.1 and similiar techniques from [34], we next derive a Putinar–Vasilescu-type Positivstellensatz for polynomial matrices with UQs.
In the homogeneous case, the result is stated as follows.
@@ 4.1. Putinar–Vasilescu-type Positivstellensatz
Suppose that Assumption 3.2 holds, $F\in{\mathbb{S}}[{\boldsymbol{x}}]^{p}$ and $G\in{\mathbb{S}}[{\boldsymbol{x}},{\boldsymbol{y}}]^{q}$ are homogeneous in ${\boldsymbol{x}}$ of even degree, and $F({\boldsymbol{x}})\succ 0$ for all ${\boldsymbol{x}}\in\mathcal{X}\setminus\{\mathbf{0}\}$. Then, there exists $N\in\mathbb{N}$ and SOS matrices $\Sigma_{0}\in{\mathbb{S}}[{\boldsymbol{x}}]^{p}$, $\Sigma\in{\mathbb{S}}[{\boldsymbol{x}},{\boldsymbol{y}}]^{pq}$ which are homogeneous in ${\boldsymbol{x}}$ and $\deg\Sigma_{0}=2N+\deg F$, $\deg_{{\boldsymbol{x}}}\Sigma=2N+\deg F-\deg_{{\boldsymbol{x}}}G$, such that
$\|{\boldsymbol{x}}\|^{2N}F({\boldsymbol{x}})=\Sigma_{0}({\boldsymbol{x}})+\int_{\mathcal{Y}}\left\langle\Sigma({\boldsymbol{x}},{\boldsymbol{y}}),G({\boldsymbol{x}},{\boldsymbol{y}})\right\rangle_{p}\mathrm{d}\nu({\boldsymbol{y}}).$
See Section 6.3 for its proof.
@@
Suppose that Assumption 3.2 holds and $F({\boldsymbol{x}})\in{\mathbb{S}}[{\boldsymbol{x}}]^{p}$ is PSD on $\mathcal{X}$. Then for any $\varepsilon>0$, there exists $N_{\varepsilon}\in\mathbb{N}$, SOS matrices $\Sigma_{0}\in{\mathbb{S}}[{\boldsymbol{x}}]^{p}$, and $\Sigma\in{\mathbb{S}}[{\boldsymbol{x}},{\boldsymbol{y}}]^{pq}$ with $\deg\Sigma_{0}\leq 2(N_{\varepsilon}+d_{F})$ and $\deg_{{\boldsymbol{x}}}\Sigma\leq 2(N_{\varepsilon}+d_{F}-d_{G})$, such that
$\theta^{N_{\varepsilon}}(F({\boldsymbol{x}})+\varepsilon\theta^{d_{F}}I_{p})=\Sigma_{0}({\boldsymbol{x}})+\int_{\mathcal{Y}}\left\langle\Sigma({\boldsymbol{x}},{\boldsymbol{y}}),G({\boldsymbol{x}},{\boldsymbol{y}})\right\rangle_{p}\mathrm{d}\nu({\boldsymbol{y}}).$
@@
    Case 2: $u_{n+1}=0$. By the definition of $d_{F}$, $x_{n+1}$ divides $\widetilde{F}(\tilde{{\boldsymbol{x}}})$. Thus, $\widetilde{F}(\tilde{{\boldsymbol{u}}})=0$. Since $\|\tilde{{\boldsymbol{u}}}\|^{2}\neq 0$, $\widetilde{F}(\tilde{{\boldsymbol{u}}})+\varepsilon\|\tilde{{\boldsymbol{u}}}\|^{2d_{F}}I_{p}\succ 0$.
Applying Theorem 4.1 to the polynomial matrices $\tilde{F}+\varepsilon\|\tilde{{\boldsymbol{x}}}\|^{2d_{F}}I_{p}$ and $\widetilde{G}$, we obtain $N_{\varepsilon}\in\mathbb{N}$ and SOS matrices $\widetilde{\Sigma}_{0}\in{\mathbb{S}}[{\boldsymbol{x}}]^{p}$, $\widetilde{\Sigma}\in{\mathbb{S}}[{\boldsymbol{x}},{\boldsymbol{y}}]^{pq}$, which are homogeneous in ${\boldsymbol{x}}$ and $\deg\widetilde{\Sigma}_{0}=2N_{\varepsilon}+2d_{F}$, $\deg_{{\boldsymbol{x}}}\widetilde{\Sigma}=2N_{\varepsilon}+2d_{F}-2d_{G}$, such that
$\|\tilde{{\boldsymbol{x}}}\|^{2N_{\varepsilon}}(\widetilde{F}(\tilde{{\boldsymbol{x}}})+\varepsilon\|\tilde{{\boldsymbol{x}}}\|^{2d_{F}}I_{p})=\widetilde{\Sigma}_{0}(\tilde{{\boldsymbol{x}}})+\int_{\mathcal{Y}}\left\langle\widetilde{\Sigma}(\tilde{{\boldsymbol{x}}},{\boldsymbol{y}}),\widetilde{G}(\tilde{{\boldsymbol{x}}},{\boldsymbol{y}})\right\rangle_{p}\mathrm{d}\nu({\boldsymbol{y}}).$ (5)
@@
4.2. Pólya-type Positivstellensatz
Pólya [38] proved that multiplying a positive homogeneous polynomial on the nonnegative orthant by some power of $\sum_{i=1}^{n}x_{i}$ yields a polynomial with nonnegative coefficients. A “positive” version of Pólya’s theorem was given by Powers and Reznick [39]. Dickinson and Povh [9] extended the result of Pólya [38] to provide a certificate for positive homogeneous polynomials on the intersection of the nonnegative orthant with a basic semialgebraic set.
We next employ Theorem 4.1 and similar techniques from [9] to present a Pólya-type Positivstellensatz for homogeneous polynomial matrices being PD on the intersection of $\mathcal{X}$ with the nonnegative orthant. Denote by ${\boldsymbol{e}}$ the column vector in ${\mathbb{R}}^{n}$ of all ones.
@@ 4.2. Pólya-type Positivstellensatz
Suppose that $F\in{\mathbb{S}}[{\boldsymbol{x}}]^{p}$ and $G\in{\mathbb{S}}[{\boldsymbol{x}},{\boldsymbol{y}}]^{q}$ are homogeneous in ${\boldsymbol{x}}$, $F({\boldsymbol{x}})\succ 0$ for all ${\boldsymbol{x}}\in{\mathbb{R}}^{n}_{+}\cap\mathcal{X}\setminus\{\mathbf{0}\}$, and $\mathcal{Y}$ is compact. Then, there exists $N\in\mathbb{N}$ and polynomial matrices $S_{0}=\sum_{|{\boldsymbol{\alpha}}|=\deg S_{0}}S_{0,{\boldsymbol{\alpha}}}{\boldsymbol{x}}^{{\boldsymbol{\alpha}}}\in{\mathbb{S}}[{\boldsymbol{x}}]^{p}$, $S=\sum_{|{\boldsymbol{\alpha}}|=\deg_{{\boldsymbol{x}}}S}S_{{\boldsymbol{\alpha}}}({\boldsymbol{y}}){\boldsymbol{x}}^{{\boldsymbol{\alpha}}}\in{\mathbb{S}}[{\boldsymbol{x}},{\boldsymbol{y}}]^{pq}$ which are homogeneous in ${\boldsymbol{x}}$ and satisfy that $S_{0,{\boldsymbol{\alpha}}}\succ 0$ for all $|{\boldsymbol{\alpha}}|=\deg S_{0}$, $S_{{\boldsymbol{\alpha}}}({\boldsymbol{y}})\succ 0$ for all $|{\boldsymbol{\alpha}}|=\deg_{{\boldsymbol{x}}}S$ and ${\boldsymbol{y}}\in\mathcal{Y}$, such that
The proof is postponed in Section 6.4.
In particular, when $\mathcal{Y}$ is a semialgebriac set defined by PMIs, the coefficient matrices $S_{{\boldsymbol{\alpha}}}({\boldsymbol{y}})$ in Theorem 4.2 have SOS-structured representations.
@@
Suppose that Assumption 3.2 holds, $F\in{\mathbb{S}}[{\boldsymbol{x}}]^{p}$ and $G\in{\mathbb{S}}[{\boldsymbol{x}},{\boldsymbol{y}}]^{q}$ are homogeneous in ${\boldsymbol{x}}$, and $F({\boldsymbol{x}})\succ 0$ for all ${\boldsymbol{x}}\in{\mathbb{R}}^{n}_{+}\cap\mathcal{X}\setminus\{\mathbf{0}\}$. Moreover, suppose that
$\mathcal{Y}=\{{\boldsymbol{y}}\in{\mathbb{R}}^{m}\mid H_{1}({\boldsymbol{y}})\succeq 0,\ldots,H_{t}({\boldsymbol{y}})\succeq 0\},\ H_{i}({\boldsymbol{y}})\in{\mathbb{S}}[{\boldsymbol{y}}]^{\ell_{i}},\ \ell_{i}\in\mathbb{N},\ i\in[t],$
and the Archimedean condition holds for the quadratic module $\mathcal{Q}(\mathbf{H})$ where $\mathbf{H}\coloneqq\{H_{1},\ldots,H_{t}\}$. Then, there exists $N\in\mathbb{N}$ and polynomial matrices $S_{0}=\sum_{|{\boldsymbol{\alpha}}|=\deg S_{0}}S_{0,{\boldsymbol{\alpha}}}{\boldsymbol{x}}^{{\boldsymbol{\alpha}}}\in{\mathbb{S}}[{\boldsymbol{x}}]^{p}$, $S=\sum_{|{\boldsymbol{\alpha}}|=\deg_{{\boldsymbol{x}}}S}S_{{\boldsymbol{\alpha}}}({\boldsymbol{y}}){\boldsymbol{x}}^{{\boldsymbol{\alpha}}}\in{\mathbb{S}}[{\boldsymbol{x}},{\boldsymbol{y}}]^{pq}$ which are homogeneous in ${\boldsymbol{x}}$ and satisfy that $S_{0,{\boldsymbol{\alpha}}}\succ 0$ for all $|{\boldsymbol{\alpha}}|=\deg S_{0}$, $S_{{\boldsymbol{\alpha}}}({\boldsymbol{y}})\in\mathcal{Q}(\mathbf{H})$ for all $|{\boldsymbol{\alpha}}|=\deg_{{\boldsymbol{x}}}S$,
@@
4.3. Lasserre-Netzer perturbative Positivstellensatz
Lasserre and Netzer [30] provided SOS approximations of polynomials that are nonnegative on $[-1,1]^{n}$ via simple high-degree perturbations. Motivated by their work, we next derive SOS-structured approximations for a polynomial matrix $F({\boldsymbol{x}})\in{\mathbb{S}}[{\boldsymbol{x}}]^{p}$ that is PSD on $\mathcal{X}\cap[-1,1]^{n}$ using high-degree perturbations.
4.3.1. Exponentially bounded matrix-valued moment problem
For a given $C>0$, Berg and Maserick [2] showed that a real sequence has representing measure supported on the hypercube $[-C,C]^{n}$ if and only if the sequence is exponentially bounded and the associated moment matrix is PSD. We next extend their results to characterize exponentially bounded matrix-valued sequence with a matrix-valued representing measure being supported on the set $\mathcal{X}\cap[-C,C]^{n}$. Beyond its intrinsic interest, we will use this result to establish a Lasserre-Netzer perturbative Positivstellensatz for polynomial matrices with UQs.
For a sequence ${\mathbf{S}}=(S_{{\boldsymbol{\alpha}}})_{{\boldsymbol{\alpha}}\in\mathbb{N}^{n}}\subseteq{\mathbb{S}}^{p}$ and $H({\boldsymbol{x}})\in{\mathbb{S}}[{\boldsymbol{x}}]^{q}$, let us recall the definition of the moment matrix $M(\mathbf{S})$ and localizing moment matrix $M(H\mathbf{S})$.
Definition 4.1.
Given a sequence $\mathbf{S}=(S_{{\boldsymbol{\alpha}}})_{{\boldsymbol{\alpha}}\in\mathbb{N}^{n}}\subseteq{\mathbb{S}}^{p}$, the associated moment matrix $M(\mathbf{S})$ is the block matrix whose block row and block column are indexed by $\mathbb{N}^{n}$ and the $({\boldsymbol{\alpha}},{\boldsymbol{\beta}})$-th block entry is $S_{{\boldsymbol{\alpha}}+{\boldsymbol{\beta}}}$ for all ${\boldsymbol{\alpha}},{\boldsymbol{\beta}}\in\mathbb{N}^{n}$. For $H\in{\mathbb{S}}[{\boldsymbol{x}}]^{q}$, the localizing matrix $M(H\mathbf{S})$ associated to $\mathbf{S}$ and $H$ is the block matrix whose block row and block column are indexed by $\mathbb{N}^{n}$ and the $({\boldsymbol{\alpha}},{\boldsymbol{\beta}})$-th block entry is $\sum_{{\boldsymbol{\gamma}}\in\mathrm{supp}(H)}S_{{\boldsymbol{\alpha}}+{\boldsymbol{\beta}}+{\boldsymbol{\gamma}}}\otimes H_{{\boldsymbol{\gamma}}}$ for all ${\boldsymbol{\alpha}},{\boldsymbol{\beta}}\in\mathbb{N}^{n}$. For $d\in\mathbb{N}$, the $d$-th order moment matrix $M_{d}(\mathbf{S})$ (resp. localizing matrix $M_{d}(H\mathbf{S})$) is the submatrix of $M(\mathbf{S})$ (resp. $M(H\mathbf{S})$) whose block row and block column are both indexed by $\mathbb{N}^{n}_{d}$.
Similar to the scalar case in [18], we give the definition of localizing moment matrix associated with $(\mathbf{S},G,\nu)$, where we write $G({\boldsymbol{x}},{\boldsymbol{y}})=\sum_{{\boldsymbol{\gamma}}\in\mbox{supp}_{{\boldsymbol{y}}}(G)}G_{{\boldsymbol{\gamma}}}({\boldsymbol{x}}){\boldsymbol{y}}^{{\boldsymbol{\gamma}}}$ with $G_{{\boldsymbol{\gamma}}}({\boldsymbol{x}})\in{\mathbb{S}}[{\boldsymbol{x}}]^{q}$.
Definition 4.2.
Given a sequence ${\mathbf{S}}=(S_{{\boldsymbol{\alpha}}})_{{\boldsymbol{\alpha}}\in\mathbb{N}^{n}}\subseteq{\mathbb{S}}^{p}$, $G({\boldsymbol{x}},{\boldsymbol{y}})\in{\mathbb{S}}[{\boldsymbol{x}},{\boldsymbol{y}}]^{q}$ and a Borel measure $\nu$ on $\mathcal{Y}$, the localizing matrix $M^{\nu}(G{\mathbf{S}})$ associated to ${\mathbf{S}}$, $G$ and $\nu$ is the block matrix whose block row and block column are indexed by $\mathbb{N}^{n}\times\mathbb{N}^{m}$ and the $(({\boldsymbol{\alpha}},{\boldsymbol{\eta}}),({\boldsymbol{\beta}},{\boldsymbol{\xi}}))$-th block entry is
$\sum_{{\boldsymbol{\gamma}}\in\mbox{supp}_{{\boldsymbol{y}}}(G)}\int_{\mathcal{Y}}{\boldsymbol{y}}^{{\boldsymbol{\gamma}}+{\boldsymbol{\eta}}+{\boldsymbol{\xi}}}\mathrm{d}\nu({\boldsymbol{y}})\left(\sum_{{\boldsymbol{\zeta}}\in\mbox{supp}_{{\boldsymbol{x}}}(G_{{\boldsymbol{\gamma}}})}S_{{\boldsymbol{\alpha}}+{\boldsymbol{\beta}}+{\boldsymbol{\zeta}}}\otimes G_{{\boldsymbol{\gamma}},{\boldsymbol{\zeta}}}\right)$
for all $({\boldsymbol{\alpha}},{\boldsymbol{\eta}}),({\boldsymbol{\beta}},{\boldsymbol{\xi}})\in\mathbb{N}^{n}\times\mathbb{N}^{m}$. For $d,k\in\mathbb{N}$, the $(d,k)$-th order localizing matrix $M^{\nu}_{d,k}(G{\mathbf{S}})$ is the submatrix of $M^{\nu}(G{\mathbf{S}})$ whose block row and block column are both indexed by $\mathbb{N}^{n}_{d}\times\mathbb{N}^{m}_{k}$.
Remark 4.1.
Throughout the paper, for a sequence indexed by $\mathbb{N}^{n}\times\mathbb{N}^{m}$ (resp., $\mathbb{N}^{n}_{d}\times\mathbb{N}^{m}_{k}$), the indices are arranged according to the order of the exponents in the monomial basis $[{\boldsymbol{x}}]_{\infty}\otimes[{\boldsymbol{y}}]_{\infty}$ (resp., $[{\boldsymbol{x}}]_{d}\otimes[{\boldsymbol{y}}]_{k}$).
With the definition of the localizing matrix $M_{d,k}^{\nu}(G{\mathbf{S}})$, we immediately have the following result.
Proposition 4.1.
Given a sequence ${\mathbf{S}}=(S_{{\boldsymbol{\alpha}}})_{{\boldsymbol{\alpha}}\in\mathbb{N}^{n}}\subseteq{\mathbb{S}}^{p}$, for any SOS matrix $\Sigma({\boldsymbol{x}},{\boldsymbol{y}})\in{\mathbb{S}}[{\boldsymbol{x}},{\boldsymbol{y}}]^{pq}$ with $\deg_{{\boldsymbol{x}}}\Sigma\leq 2d$ and $\deg_{{\boldsymbol{y}}}\Sigma\leq 2k$,

### v4

@@ 4. Positivstellensätze for polynomial matrices with UQs and without the Archimedean condition
The results in Section 3 are derived under the Archimedean condition on the quadratic module generated by $(G,\nu)$, which requires $\mathcal{X}$ to be bounded. Many efforts have been made to provide SOS-structured representations for polynomials that are PD on a semialgebraic set without the Archimedean condition. The goal of this section is to extend some of these well-known results to the setting of polynomial matrices with UQs.
4.1. A Positivstellensatz for the non-compact case
Let $\theta\coloneqq 1+\|{\boldsymbol{x}}\|^{2}$. For every nonnegative polynomial $f\in{\mathbb{R}}[{\boldsymbol{x}}]$ on a general basic semialgebraic set, Putinar and Vasilescu [38] proved that for a given $\varepsilon>0$ and $2d\geq\deg f$, there exists a nonnegative integer $k$ such that $\theta^{k}(f+\varepsilon\theta^{d})$ belongs to the quadratic module associated with the basic semialgebraic set. Using Jacobi’s technique [18], Mai et al. [30] provided an alternative proof of Putinar and Vasilescu’s result with an effective degree bound on polynomials involved in such certificates.
Building on Theorem 3.1 and similar techniques from [30], we next derive matrix-valued Positivstellensätz for the non-compact case, incorporating universal quantifiers. In the homogeneous case, the result is stated as follows. See Section 6.2 for its proof.
@@ 4.1. A Positivstellensatz for the non-compact case
Suppose that Assumption 3.2 holds, $F\in{\mathbb{S}}[{\boldsymbol{x}}]^{p}$ and $G\in{\mathbb{S}}[{\boldsymbol{x}},{\boldsymbol{y}}]^{q}$ are homogeneous in $\boldsymbol{x}$} of even degree, and $F({\boldsymbol{x}})\succ 0$ for all ${\boldsymbol{x}}\in\mathcal{X}\setminus\{\mathbf{0}\}$. Then, there exists $N\in\mathbb{N}$ and SOS matrices $_{0}\in{\mathbb{S}}[{\boldsymbol{x}}]^{p}, S𝕊$$\in{\mathbb{S}}[{\boldsymbol{x}},{\boldsymbol{y}}]^{pq}$ that are homogeneous in $\boldsymbol{x}$} and $\deg S_{0}=2N+\deg F$, $\deg_{{\boldsymbol{x}}}S=2N+\deg F-\deg_{{\boldsymbol{x}}}G$, such that
$\|{\boldsymbol{x}}\|^{2N}F({\boldsymbol{x}})=S_{0}({\boldsymbol{x}})+\int_{\mathcal{Y}}\left\langle S({\boldsymbol{x}},{\boldsymbol{y}}),G({\boldsymbol{x}},{\boldsymbol{y}})\right\rangle_{p}\mathrm{d}\nu({\boldsymbol{y}}).$
@@
Suppose that Assumption 3.2 holds and $F({\boldsymbol{x}})\in{\mathbb{S}}[{\boldsymbol{x}}]^{p}$ is PSD on $\mathcal{X}$. Then for any $\varepsilon>0$, there exists $N_{\varepsilon}\in\mathbb{N}$, SOS matrices $S_{0}\in{\mathbb{S}}[{\boldsymbol{x}}]^{p}$, and $S\in{\mathbb{S}}[{\boldsymbol{x}},{\boldsymbol{y}}]^{pq}$ with $\deg S_{0}\leq 2(N_{\varepsilon}+d_{F})$ and $\deg_{{\boldsymbol{x}}}S\leq 2(N_{\varepsilon}+d_{F}-d_{G})$, such that
$\theta^{N_{\varepsilon}}(F({\boldsymbol{x}})+\varepsilon\theta^{d_{F}}I_{p})=S_{0}({\boldsymbol{x}})+\int_{\mathcal{Y}}\left\langle S({\boldsymbol{x}},{\boldsymbol{y}}),G({\boldsymbol{x}},{\boldsymbol{y}})\right\rangle_{p}\mathrm{d}\nu({\boldsymbol{y}}).$
@@
[…]
Applying Theorem 4.1 to the polynomial matrices $\tilde{F}+\varepsilon\|\tilde{{\boldsymbol{x}}}\|^{2d_{F}}I_{p}$ and $\widetilde{G}$, we obtain $N_{\varepsilon}\in\mathbb{N}$ and SOS matrices $\widetilde{S}_{0}\in{\mathbb{S}}[{\boldsymbol{x}}]^{p}$, $\widetilde{S}\in{\mathbb{S}}[{\boldsymbol{x}},{\boldsymbol{y}}]^{pq}$, which are homogeneous in $\boldsymbol{x}$} and $\deg\widetilde{S}_{0}=2N_{\varepsilon}+2d_{F}$, $\deg_{{\boldsymbol{x}}}\widetilde{S}=2N_{\varepsilon}+2d_{F}-2d_{G}$, such that
$\|\tilde{{\boldsymbol{x}}}\|^{2N_{\varepsilon}}(\widetilde{F}(\tilde{{\boldsymbol{x}}})+\varepsilon\|\tilde{{\boldsymbol{x}}}\|^{2d_{F}}I_{p})=\widetilde{S}_{0}(\tilde{{\boldsymbol{x}}})+\int_{\mathcal{Y}}\left\langle\widetilde{S}(\tilde{{\boldsymbol{x}}},{\boldsymbol{y}}),\widetilde{G}(\tilde{{\boldsymbol{x}}},{\boldsymbol{y}})\right\rangle_{p}\mathrm{d}\nu({\boldsymbol{y}}).$ (6)
@@
4.2. A Positivstellensatz on the nonnegative orthant
Pólya [35] proved that multiplying a positive homogeneous polynomial on the nonnegative orthant by some power of $\sum_{i=1}^{n}x_{i}$ yields a polynomial with nonnegative coefficients. Powers and Reznick [36] gave a ”positive” version of Pólya’s theorem. Dickinson and Povh [7] extended the result of Pólya [35] to provide a certificate for positive homogeneous polynomials on the intersection of the nonnegative orthant with a basic semialgebraic set.
We next employ Theorem 4.1 and similar techniques from [7] to present a Positivstellensatz for homogeneous polynomial matrices being PD on the intersection of $\mathcal{X}$ with the nonnegative orthant. Denote by $\boldsymbol{e}$} the column vector in ${\mathbb{R}}^{n}$ of all ones.
@@ 4.2. A Positivstellensatz on the nonnegative orthant
Suppose that $F\in{\mathbb{S}}[{\boldsymbol{x}}]^{p}$ and $G\in{\mathbb{S}}[{\boldsymbol{x}},{\boldsymbol{y}}]^{q}$ are homogeneous in $\boldsymbol{x}$}, $F({\boldsymbol{x}})\succ 0$ for all ${\boldsymbol{x}}\in{\mathbb{R}}^{n}_{+}\cap\mathcal{X}\setminus\{\mathbf{0}\}$, and $\mathcal{Y}$ is compact. Then, there exists $N\in\mathbb{N}$ and polynomial matrices
$P_{0}=\sum_{|{\boldsymbol{\alpha}}|=N+\deg F}P_{0,{\boldsymbol{\alpha}}}{\boldsymbol{x}}^{{\boldsymbol{\alpha}}}\in{\mathbb{S}}[{\boldsymbol{x}}]^{p},\quad P=\sum_{|{\boldsymbol{\alpha}}|=N+\deg F-\deg_{{\boldsymbol{x}}}G}P_{{\boldsymbol{\alpha}}}({\boldsymbol{y}}){\boldsymbol{x}}^{{\boldsymbol{\alpha}}}\in{\mathbb{S}}[{\boldsymbol{x}},{\boldsymbol{y}}]^{pq}$
which are homogeneous in $\boldsymbol{x}$} and satisfy $P_{0,{\boldsymbol{\alpha}}}\succ 0$ for all $|{\boldsymbol{\alpha}}|=N+\deg F$, $P_{{\boldsymbol{\alpha}}}({\boldsymbol{y}})\succ 0$ for all $|{\boldsymbol{\alpha}}|=N+\deg F-\deg_{{\boldsymbol{x}}}G$ and ${\boldsymbol{y}}\in\mathcal{Y}$, such that
$({\boldsymbol{e}}^{\intercal}{\boldsymbol{x}})^{N}F({\boldsymbol{x}})=P_{0}({\boldsymbol{x}})+\int_{\mathcal{Y}}\left\langle P({\boldsymbol{x}},{\boldsymbol{y}}),G({\boldsymbol{x}},{\boldsymbol{y}})\right\rangle_{p}\mathrm{d}\nu({\boldsymbol{y}}).$
The proof is postponed to Section 6.3.
In particular, when $\mathcal{Y}$ is a semialgebriac set defined by PMIs, the coefficient matrices $P_{{\boldsymbol{\alpha}}}({\boldsymbol{y}})$ in Theorem 4.2 have SOS-structured representations.
@@
Suppose that $F\in{\mathbb{S}}[{\boldsymbol{x}}]^{p}$ and $G\in{\mathbb{S}}[{\boldsymbol{x}},{\boldsymbol{y}}]^{q}$ are homogeneous in $\boldsymbol{x}$}, and $F({\boldsymbol{x}})\succ 0$ for all ${\boldsymbol{x}}\in{\mathbb{R}}^{n}_{+}\cap\mathcal{X}\setminus\{\mathbf{0}\}$. Moreover, suppose that
$\mathcal{Y}=\{{\boldsymbol{y}}\in{\mathbb{R}}^{m}\mid H_{1}({\boldsymbol{y}})\succeq 0,\ldots,H_{t}({\boldsymbol{y}})\succeq 0\},\ H_{i}({\boldsymbol{y}})\in{\mathbb{S}}[{\boldsymbol{y}}]^{\ell_{i}},\ \ell_{i}\in\mathbb{N},\ i\in[t],$ (7)
and the Archimedean condition holds for the quadratic module $\mathcal{Q}^{pq}(\mathbf{H})$ where $\mathbf{H}\coloneqq\{H_{1},\ldots,H_{t}\}$. Then, there exists $N\in\mathbb{N}$ and polynomial matrices
$P_{0}=\sum_{|{\boldsymbol{\alpha}}|=N+\deg F}P_{0,{\boldsymbol{\alpha}}}{\boldsymbol{x}}^{{\boldsymbol{\alpha}}}\in{\mathbb{S}}[{\boldsymbol{x}}]^{p},\quad P=\sum_{|{\boldsymbol{\alpha}}|=N+\deg F-\deg_{{\boldsymbol{x}}}G}P_{{\boldsymbol{\alpha}}}({\boldsymbol{y}}){\boldsymbol{x}}^{{\boldsymbol{\alpha}}}\in{\mathbb{S}}[{\boldsymbol{x}},{\boldsymbol{y}}]^{pq}$
which are homogeneous in $\boldsymbol{x}$} and satisfy $P_{0,{\boldsymbol{\alpha}}}\succ 0$ for all $|{\boldsymbol{\alpha}}|=N+\deg F$, $P_{{\boldsymbol{\alpha}}}({\boldsymbol{y}})\in\mathcal{Q}^{pq}(\mathbf{H})$ for all $|{\boldsymbol{\alpha}}|=N+\deg F-\deg_{{\boldsymbol{x}}}G$, such that
$({\boldsymbol{e}}^{\intercal}{\boldsymbol{x}})^{N}F({\boldsymbol{x}})=P_{0}({\boldsymbol{x}})+\int_{\mathcal{Y}}\left\langle P({\boldsymbol{x}},{\boldsymbol{y}}),G({\boldsymbol{x}},{\boldsymbol{y}})\right\rangle_{p}\mathrm{d}\nu({\boldsymbol{y}}).$
@@
In analogy with Corollary 4.1, we can obtain the following inhomogeneous counterpart of Theorem 4.2. We omit the similar proof for the sake of clarity.
Corollary 4.3.
Suppose that $F\in{\mathbb{S}}[{\boldsymbol{x}}]^{p}$ and $G\in{\mathbb{S}}[{\boldsymbol{x}},{\boldsymbol{y}}]^{q}$, $F({\boldsymbol{x}})$ is PSD on ${\mathbb{R}}^{n}_{+}\cap\mathcal{X}$, and $\mathcal{Y}$ is compact. Then for any $\varepsilon>0$, there exists $N_{\varepsilon}\in\mathbb{N}$ and polynomial matrices
$P_{0}=\sum_{|{\boldsymbol{\alpha}}|\leq N_{\varepsilon}+2d_{F}}P_{0,{\boldsymbol{\alpha}}}{\boldsymbol{x}}^{{\boldsymbol{\alpha}}}\in{\mathbb{S}}[{\boldsymbol{x}}]^{p},\quad P=\sum_{|{\boldsymbol{\alpha}}|\leq N_{\varepsilon}+2d_{F}-2d_{G}}P_{{\boldsymbol{\alpha}}}({\boldsymbol{y}}){\boldsymbol{x}}^{{\boldsymbol{\alpha}}}\in{\mathbb{S}}[{\boldsymbol{x}},{\boldsymbol{y}}]^{pq}$
satisfying $P_{0,{\boldsymbol{\alpha}}}\succ 0$ for all $|{\boldsymbol{\alpha}}|\leq N_{\varepsilon}+2d_{F}$, $P_{{\boldsymbol{\alpha}}}({\boldsymbol{y}})\succ 0$ for all $|{\boldsymbol{\alpha}}|\leq N_{\varepsilon}+2d_{F}-2d_{G}$ and ${\boldsymbol{y}}\in\mathcal{Y}$, and